A baseball team scored the following number of runs in its games this season:

6, 2, 5, 9, 11, 4, 5, 8, 6, 7, 5

There is one more game in the season. If the team wants to end the season with an average of at least 6 runs pergame, what is the least number of runs the team must score in the final game of the season?
Answer #1

Let the team need to score at least x runs to end up with an average of at least 6.

Taking the minimum average score which would be 6, sum = 6*12 = 72

Sum of given runs = 6 + 2 + 5 + 9 + 11 + 4 + 5 + 8 + 6 + 7 + 5 = 68

Hence,

68 + x = 72

x = 72 - 68

x = 4

So the team need to score atleast 4 runs in the last game to end up with an average of atleast 6 runs.